# REP: Resource-Efficient Prompting for Rehearsal-Free Continual Learning

### v4

@@ 1 Introduction
Continual learning (CL) trains neural network models on multiple sequential tasks, where each task may include data distributions that diverge from previously encountered data. A crucial challenge for any CL algorithm is to effectively address catastrophic forgetting [23]. Severe forgetting occurs when a model rapidly loses previously learned knowledge while adapting to new tasks, significantly affecting the model’s reliability and accuracy on earlier tasks.
Moreover, many of today’s AI services are designed for on-device scenarios to securely learn tasks locally [19, 10, 42]. In on-device CL, improving computational efficiency is crucial, as it directly reduces energy usage and enhances the durability of edge devices. Meanwhile, since device memory capacity often acts as a hard constraint, CL tasks that exhaust all available memory can cause system crashes due to out-of-memory errors [9]. Given typically limited memory sizes (1–8GB), achieving high memory efficiency is essential to enable diverse on-device deployment scenarios in practice.
In this paper, we propose REP (Fig.˜1), a framework for resource efficiency in prompt-based rehearsal-free CL methods on frozen, pre-trained vision transformers (ViTs) [37, 38, 35, 31, 14, 13, 30, 33]. Prompts are a small set of parameters that progressively learn incoming tasks to combat forgetting. Updates to these compact prompts incur minimum data writes without significantly harming the lifespan of device storage, which typically sustains up to 10K of writes per location. This makes prompt-based rehearsal-free methods well-suited for on-device continual learning.
@@
REP is built upon our analysis of cost-accuracy trade-offs throughout the end-to-end learning process, with two key design insights. (1) The prompt selection stage, which constructs a prompt subset to augment input or intermediate data is highly amenable to numerous promising options with fast approximations. (2) In contrast, the prompt update stage, which involves forward-backward passes over the backbone model, presents a range of optimizations with vastly different cost-accuracy trade-offs. Based on these insights, we develop three complementary techniques for both stages that trade negligible accuracy drop for high resource efficiency: (1) random projection-based lightweight surrogate model for prompt selection, and (2) adaptive token merging (AToM) and (3) adaptive layer dropping (ALD) for prompt update. Notably, AToM and ALD non-uniformly skip parts of the data and model layers to reduce training costs while preserving critical task-specific knowledge. These approaches are inspired by our analysis (Fig.˜2) and prior works [37, 28] demonstrating that shallower layers in pre-trained ViTs are more sensitive to new tasks than deeper layers.
@@ 2 Related work
Class-incremental learning.  Our work focuses on a class-incremental learning (CIL) setting, where each task introduces new classes within the same domain, but task IDs are unavailable during inference [8]. In CNN-based approaches, rehearsal-based methods that replay stored samples remain dominant due to their strong performance [1, 4, 5, 26, 29]. Recent methods such as BudgetCL [27] and CarM [20] adopt data-driven strategies that outperform optimization-based methods under computational constraints. MEMO [43] trades stored samples for task-specific layers when beneficial. We compare our prompt-based approach with these methods and show superior performance under resource constraints.
Continual learning for the edge.  Recent efforts in on-device learning have emphasized memory and energy efficiency, primarily outside the CL domain [9, 34, 36]. A few studies have extended CL to edge devices: Hayes and Kanan [10] investigated online CL with CNNs for embedded systems; Kwon et al. [19] analyzed rehearsal vs. regularization methods in terms of storage, compute, and accuracy trade-offs; and Ma et al. [22] introduced Miro, a platform that dynamically configures CNN-based CL to reduce energy consumption within memory constraints. However, none of these works address the challenge of resource-efficient vision transformers (ViTs) for on-device CL, which is the primary focus of our work.
Prompting for continual learning.  Prompting, which provides explicit instructions or queries to the model, has proven effective for ViTs in CL by enabling adaptation to new tasks while preserving prior knowledge [13, 14, 30, 31, 33, 35, 37, 38]. L2P [38] and DualPrompt [37] retrieve task-relevant prompts from a shared pool using prompt tuning and prefix tuning, respectively. CODA-Prompt [31] improves prompting capacity with attention-conditioned prompts, while OVOR [13] uses a single prompt to accelerate selection. HiDe-Prompt [33] extends supervised prompt-based CL to self-supervised settings through hierarchical optimization. ConvPrompt [30] employs convolution-based, layer-wise prompts to improve knowledge transfer with negligible overhead. LAE [7] interprets prompts as a form of parameter-efficient fine-tuning, incorporating modules such as adapters or LoRA. As shown in Section˜4, our techniques can be applied to these methods to enhance resource efficiency without significant accuracy loss.
@@ 3.1 Motivation: challenges in prompt selection and update
where $f_{\text{update}}$ denotes the main backbone, and $\epsilon_{1}$ and $\epsilon_{2}$ control the influence of the prompt-specific and auxiliary losses. Different methods instantiate this loss differently. For example, [37, 38] omit the auxiliary loss by setting $\epsilon_{1}=1,\epsilon_{2}=0$, whereas [31, 33] tune both coefficients.
@@ 3.2.1 Guiding prompt selection with a lightweight surrogate model
To reduce the computational cost of prompt selection, we use a lightweight surrogate model $f_{\text{efficient}}$ with reduced depth and width. Instead of reusing a large backbone such as ViT-L, we adopt a compact pre-trained ViT-Ti model, which effectively captures essential representations. Given an input $x_{i}^{j}$, $f_{\text{efficient}}$ produces a low-dimensional query feature $q_{\text{efficient}}(x_{i}^{j})\in\mathbb{R}^{d}$ (where $d\leq D$). We then apply a fixed, non-trainable random projection $\phi$ [24] to map this feature back to the original $D$-dimensional space for prompt selection:
@@
Empirically, this strategy preserves around 97% of the representational similarity to a ViT-L-based query model, as measured by centered kernel alignment [17]. Despite residing in a low-dimensional space, $p^{*}_{\text{efficient}}$ effectively approximates the large model’s representations, substantially reducing computational cost (see Table˜2).
@@ 3.2.2 Analysis of the frozen backbone during prompt update
A key insight of REP is that not all frozen blocks contribute equally to the final loss $L$. To validate this, we analyze the feature representations of pre-trained transformer blocks with prompts using mean attention distance, a metric previously used to study layer-wise representations in ViT models during pre-training [6, 28].
For a given attention head within a frozen block, let $(x_{q},y_{q})$ denote the position of a query patch, and $(x_{i},y_{i})$ the positions of the patches the head attends to, with corresponding attention weights $a_{i}$. The distance $d_{i}$ between patches—typically defined as Euclidean or pixel distance—is given by $d_{i}=(x_{i}-x_{q})^{2}+(y_{i}-y_{q})^{2}$. We then compute the weighted mean distance as $\frac{\sum_{i}a_{i}\cdot d_{i}}{\sum_{i}a_{i}}$. Following [6, 28], a smaller distance indicates attention focused on a narrow region (local information), while a larger distance suggests more distributed attention across the input (global information). We measure the average mean attention distance of L2P with the ViT-L backbone along two dimensions: layer ID and attention head, during adaptation to a new task in Split ImageNet-R (10 tasks).
As shown in Fig.˜2(a), shallower layers exhibit a wider range of attention distances and tend to focus on both localized and global regions. In contrast, deeper layers consistently produce higher values, reflecting a shift toward more global contextual information. This pattern aligns well with prior findings in [6, 28]. A closer look at a few selected layers in Fig.˜2(b) further reveals that attention distances across heads vary more widely in shallow layers than in deeper ones, confirming that shallower layers capture more important representations of the input.
@@ 3.2.3 Adaptive token merging (AToM)
We explain how we bring the above insights into practice through two compute-skipping techniques. We first consider the data-efficient compute-skipping method via token merging. Conventional token merging (ToMe) [2] reduces the number of tokens by uniformly merging $n$ redundant tokens per layer, controlled by a fixed scheduler $r$. The scheduler function $r(l)\rightarrow n$ is applied to each layer $l$, and according to [2], it merges all tokens, including the prompts added by the token selection process. However, insights from Fig.˜2 and our additional analysis highlight two major problems.
@@ 3.2.4 Adaptive layer dropping (ALD)
Inspired by insights from Fig.˜2 and prior work on progressive layer dropping (PLD) [41] in the NLP domain, we propose adaptive layer dropping (ALD; Algorithm˜2) with two key features: (1) the dropping schedule considers both temporal and spatial dimensions, and (2) it manages to drop layers non-uniformly to preserve critical representation features in shallower layers. [^1] The spatial dimension refers to model layers, each processing input features at different levels of abstraction. On the contrary, PLD considers only the temporal aspect and does not differentiate between layers when dropping. This results in poorer performance compared to ALD, as shown in Table˜3.
@@ 3.2.5 Discussion
Compact backbone networks (e.g., CNNs).  Although CNNs are generally more compact than vision transformers (ViTs), they consistently underperform in the CL setting, particularly under tight resource constraints. Our preliminary studies in Appendix˜D demonstrate that state-of-the-art CNN-based CL methods, such as BudgetCL [27] and MEMO [43], achieve significantly lower accuracy compared to ViT-based approaches.
@@ 4 Experiments
We focus on the popular disjoint-CIL setup, where each task comprises a distinct set of non-overlapping classes and samples from old classes are not given in future tasks [29, 4, 8]. In this section, we demonstrate that (1) REP enhances resource efficiency in prompt-based rehearsal-free CL across various backbone model sizes and (2) its core components, ALD and AToM, are applicable even when CL methods do not employ prompting.
Data generation.  To organize task streams, we use three image classification datasets: CIFAR-100 (100 classes) [18], ImageNet-R (200 classes) [12], and PlantDisease (38 classes) [25]. Out of these datasets, ImageNet-R is known for exhibiting much higher intra-class variability and an uneven class-size distribution among images. We divide CIFAR-100 and ImageNet-R into 10 tasks to create Split CIFAR-100 (i.e., 10 classes per task) and Split ImageNet-R (i.e., 20 classes per task), respectively [38, 37, 31]. For PlantDisease, we drop 3 plant disease classes with very few images and organize the remaining 35 classes into 7 tasks to create Split PlantDisease (i.e., 5 classes per task). We also present additional results with varying task lengths (e.g., 5 tasks and 20 tasks), and an additional Split CUB-200 dataset [32] in Appendix˜F and Appendix˜G, respectively.
Methods.  We employ L2P [38], DualPrompt [37], CODA-Prompt [31], HiDe-Prompt [33], ConvPrompt [30], LAE [7], and OVOR [13] as representative prompting methods. These methods capitalize on ImageNet pre-trained models as backbones: ViT-L (307M), ViT-B (86M), and ViT-Ti (5.8M)—ViT-Ti is one of the smallest vision transformer models to our knowledge. Detailed settings for all methods are provided in Appendix˜B.
Hardware and metrics.  We use an NVIDIA RTX 3090 to cover a wide range of memory capacities while maintaining consistent computational power. Specifically, we limit the GPU memory to emulate resource-constrained devices, ranging from 1GB (compatible with Raspberry Pi) to 16GB (compatible with NVIDIA Jetson AGX Xavier). For computational efficiency, we measure iteration time, which is linearly correlated with GPU energy usage [22]. We adopt the final average accuracy (higher is better) and forgetting (lower is better) from various previous works [37, 38, 31, 33]. In Appendix˜M, we further evaluate REP directly on NVIDIA Jetson TX2 with 8GB of memory, although it is applicable beyond this specific device.
@@ 4.2 Ablation studies
Table 3: Effect of using (1) conventional acceleration methods, (2) intermediate layer-based prompt selection method, and (3) Random Drop with diverse drop ratio, compared to our proposed methods.
@@
Algorithm validation.  Conventional token merging (ToMe) [2] and layer dropping (PLD) [41] are specifically designed to accelerate the training of transformer-based models across various domains. To validate the importance of incorporating our adaptive techniques instead in CL, we first evaluate how REP performs in case AToM and ALD are replaced with ToMe and PLD, respectively. The results are presented in (1) of Table˜3. Although applying ToMe or PLD improves resource efficiency over the native L2P baseline in Table˜1, it results in an accuracy drop of 5.1% or 2.0%, respectively, compared to using our techniques.
@@
Comparison with FIM-based token merging.  We also investigate whether additional gradient information, e.g. through the Fisher Information Matrix (FIM) [16], could serve as an alternative to AToM. Specifically, we compare AToM with a FIM-based, gradient-informed merging strategy using L2P with ViT-L on Split ImageNet-R (10 tasks), as shown in (3) of Table˜3. Although this approach yields modest accuracy gains (0.1–0.2%), it incurs roughly a 10% increase in training time and memory usage. By contrast, AToM preserves prompt gradients without such overhead, offering a more balanced trade-off between efficiency and forgetting mitigation. This result does not point to a fundamental limitation of FIM-based merging, but rather underscores the difficulty of achieving practical efficiency improvements in its current form.
@@
AToM and ALD intensity.  Table˜4 shows the effects of varying intensities of AToM and ALD, focusing on the number of merged tokens ($n$) in AToM and the keep ratio ($\theta$) in ALD. AToM appears to maintain stable accuracy trends even as more tokens are merged. In contrast, in the case of ALD, a lower keep ratio appears to improve resource efficiency by reducing training time. However, it can markedly impair model accuracy if the ratio is too low. Overall, when used with carefully selected hyperparameters, AToM and ALD can effectively balance resource efficiency and accuracy. Notably, forgetting remains stable when using the default settings ($n$ = 8 and $\theta$ = 0.5). In Appendix˜E, we extend the above study to two additional prompt-based CL methods, including HiDe-Prompt [33] and ConvPrompt [30].

### v6

@@ 1 Introduction
Continual learning (CL) trains neural network models on multiple sequential tasks, where each task may include data distributions that diverge from previously encountered data. A crucial challenge for any CL algorithm is to effectively address catastrophic forgetting [mccloskeyC89]. Severe forgetting occurs when a model rapidly loses previously learned knowledge while adapting to new tasks, significantly affecting the model’s reliability and accuracy on earlier tasks.
Moreover, many of today’s AI services are designed for on-device scenarios to securely learn tasks locally [kwon2021exploring, hayes2022online, e-domainil]. In on-device CL, improving computational efficiency is crucial, as it directly reduces energy usage and enhances the durability of edge devices. Meanwhile, since device memory capacity often acts as a hard constraint, CL tasks that exhaust all available memory can cause system crashes due to out-of-memory errors [sage]. Given typically limited memory sizes (1–8GB), achieving high memory efficiency is essential to enable diverse on-device deployment scenarios in practice.
In this paper, we propose REP (Fig.˜1), a framework for resource efficiency in prompt-based rehearsal-free CL methods on frozen, pre-trained vision transformers (ViTs) [dualp, l2p, sprompts, codap, dap, ovor, convp, hidep]. Prompts are a small set of parameters that progressively learn incoming tasks to combat forgetting. Updates to these compact prompts incur minimum data writes without significantly harming the lifespan of device storage, which typically sustains up to 10K of writes per location. This makes prompt-based rehearsal-free methods well-suited for on-device continual learning.
@@
REP is built upon our analysis of cost-accuracy trade-offs throughout the end-to-end learning process, with two key design insights. (1) The prompt selection stage, which constructs a prompt subset to augment input or intermediate data is highly amenable to numerous promising options with fast approximations. (2) In contrast, the prompt update stage, which involves forward-backward passes over the backbone model, presents a range of optimizations with vastly different cost-accuracy trade-offs. Based on these insights, we develop three complementary techniques for both stages that trade negligible accuracy drop for high resource efficiency: (1) random projection-based lightweight surrogate model for prompt selection, and (2) adaptive token merging (AToM) and (3) adaptive layer dropping (ALD) for prompt update. Notably, AToM and ALD non-uniformly skip parts of the data and model layers to reduce training costs while preserving critical task-specific knowledge. These approaches are inspired by our analysis (Fig.˜2) and prior works [dualp, vitnature] demonstrating that shallower layers in pre-trained ViTs are more sensitive to new tasks than deeper layers.
@@ 2 Related work
Class-incremental learning.  Our work focuses on a class-incremental learning (CIL) setting, where each task introduces new classes within the same domain, but task IDs are unavailable during inference [gepperthH16]. In CNN-based approaches, rehearsal-based methods that replay stored samples remain dominant due to their strong performance [rainbow, castro2018eccv, rwalk, gdumb, icarl]. Recent methods such as BudgetCL [budgetcl] and CarM [carm] adopt data-driven strategies that outperform optimization-based methods under computational constraints. MEMO [memo] trades stored samples for task-specific layers when beneficial. We compare our prompt-based approach with these methods and show superior performance under resource constraints.
Continual learning for the edge.  Recent efforts in on-device learning have emphasized memory and energy efficiency, primarily outside the CL domain [sage, melon, e2train]. A few studies have extended CL to edge devices: hayes2022online investigated online CL with CNNs for embedded systems; kwon2021exploring analyzed rehearsal vs. regularization methods in terms of storage, compute, and accuracy trade-offs; and miro introduced Miro, a platform that dynamically configures CNN-based CL to reduce energy consumption within memory constraints. However, none of these works address the challenge of resource-efficient vision transformers (ViTs) for on-device CL, which is the primary focus of our work.
Prompting for continual learning.  Prompting, which provides explicit instructions or queries to the model, has proven effective for ViTs in CL by enabling adaptation to new tasks while preserving prior knowledge [ovor, dap, convp, codap, hidep, sprompts, dualp, l2p]. L2P [l2p] and DualPrompt [dualp] retrieve task-relevant prompts from a shared pool using prompt tuning and prefix tuning, respectively. CODA-Prompt [codap] improves prompting capacity with attention-conditioned prompts, while OVOR [ovor] uses a single prompt to accelerate selection. HiDe-Prompt [hidep] extends supervised prompt-based CL to self-supervised settings through hierarchical optimization. ConvPrompt [convp] employs convolution-based, layer-wise prompts to improve knowledge transfer with negligible overhead. LAE [lae] interprets prompts as a form of parameter-efficient fine-tuning, incorporating modules such as adapters or LoRA. As shown in Section˜4, our techniques can be applied to these methods to enhance resource efficiency without significant accuracy loss.
@@ 3.1 Motivation: challenges in prompt selection and update
where $f_{\text{update}}$ denotes the main backbone, and $\epsilon_{1}$ and $\epsilon_{2}$ control the influence of the prompt-specific and auxiliary losses. Different methods instantiate this loss differently. For example, [dualp, l2p] omit the auxiliary loss by setting $\epsilon_{1}=1,\epsilon_{2}=0$, whereas [codap, hidep] tune both coefficients.
@@ 3.2.1 Guiding prompt selection with a lightweight surrogate model
To reduce the computational cost of prompt selection, we use a lightweight surrogate model $f_{\text{efficient}}$ with reduced depth and width. Instead of reusing a large backbone such as ViT-L, we adopt a compact pre-trained ViT-Ti model, which effectively captures essential representations. Given an input $x_{i}^{j}$, $f_{\text{efficient}}$ produces a low-dimensional query feature $q_{\text{efficient}}(x_{i}^{j})\in\mathbb{R}^{d}$ (where $d\leq D$). We then apply a fixed, non-trainable random projection $\phi$ [ranpac] to map this feature back to the original $D$-dimensional space for prompt selection:
@@
Empirically, this strategy preserves around 97% of the representational similarity to a ViT-L-based query model, as measured by centered kernel alignment [cka]. Despite residing in a low-dimensional space, $p^{*}_{\text{efficient}}$ effectively approximates the large model’s representations, substantially reducing computational cost (see Table˜2).
@@ 3.2.2 Analysis of the frozen backbone during prompt update
A key insight of REP is that not all frozen blocks contribute equally to the final loss $L$. To validate this, we analyze the feature representations of pre-trained transformer blocks with prompts using mean attention distance, a metric previously used to study layer-wise representations in ViT models during pre-training [vit_original, vitnature].
For a given attention head within a frozen block, let $(x_{q},y_{q})$ denote the position of a query patch, and $(x_{i},y_{i})$ the positions of the patches the head attends to, with corresponding attention weights $a_{i}$. The distance $d_{i}$ between patches—typically defined as Euclidean or pixel distance—is given by $d_{i}=(x_{i}-x_{q})^{2}+(y_{i}-y_{q})^{2}$. We then compute the weighted mean distance as $\frac{\sum_{i}a_{i}\cdot d_{i}}{\sum_{i}a_{i}}$. Following [vit_original, vitnature], a smaller distance indicates attention focused on a narrow region (local information), while a larger distance suggests more distributed attention across the input (global information). We measure the average mean attention distance of L2P with the ViT-L backbone along two dimensions: layer ID and attention head, during adaptation to a new task in Split ImageNet-R (10 tasks).
As shown in Fig.˜2(a), shallower layers exhibit a wider range of attention distances and tend to focus on both localized and global regions. In contrast, deeper layers consistently produce higher values, reflecting a shift toward more global contextual information. This pattern aligns well with prior findings in [vit_original, vitnature]. A closer look at a few selected layers in Fig.˜2(b) further reveals that attention distances across heads vary more widely in shallow layers than in deeper ones, confirming that shallower layers capture more important representations of the input.
@@ 3.2.3 Adaptive token merging (AToM)
We explain how we bring the above insights into practice through two compute-skipping techniques. We first consider the data-efficient compute-skipping method via token merging. Conventional token merging (ToMe) [token_merging] reduces the number of tokens by uniformly merging $n$ redundant tokens per layer, controlled by a fixed scheduler $r$. The scheduler function $r(l)\rightarrow n$ is applied to each layer $l$, and according to [token_merging], it merges all tokens, including the prompts added by the token selection process. However, insights from Fig.˜2 and our additional analysis highlight two major problems.
@@ 3.2.4 Adaptive layer dropping (ALD)
Inspired by insights from Fig.˜2 and prior work on progressive layer dropping (PLD) [progressive_drop] in the NLP domain, we propose adaptive layer dropping (ALD; Algorithm˜2) with two key features: (1) the dropping schedule considers both temporal and spatial dimensions, and (2) it manages to drop layers non-uniformly to preserve critical representation features in shallower layers. [^1] The spatial dimension refers to model layers, each processing input features at different levels of abstraction. On the contrary, PLD considers only the temporal aspect and does not differentiate between layers when dropping. This results in poorer performance compared to ALD, as shown in Table˜3.
@@ 3.2.5 Discussion
Compact backbone networks (e.g., CNNs).  Although CNNs are generally more compact than vision transformers (ViTs), they consistently underperform in the CL setting, particularly under tight resource constraints. Our preliminary studies in Appendix˜D demonstrate that state-of-the-art CNN-based CL methods, such as BudgetCL [budgetcl] and MEMO [memo], achieve significantly lower accuracy compared to ViT-based approaches.
@@ 4 Experiments
We focus on the popular disjoint-CIL setup, where each task comprises a distinct set of non-overlapping classes and samples from old classes are not given in future tasks [icarl, castro2018eccv, gepperthH16]. In this section, we demonstrate that (1) REP enhances resource efficiency in prompt-based rehearsal-free CL across various backbone model sizes and (2) its core components, ALD and AToM, are applicable even when CL methods do not employ prompting.
Data generation.  To organize task streams, we use three image classification datasets: CIFAR-100 (100 classes) [cifar], ImageNet-R (200 classes) [imagenet_r], and PlantDisease (38 classes) [plantdisease]. Out of these datasets, ImageNet-R is known for exhibiting much higher intra-class variability and an uneven class-size distribution among images. We divide CIFAR-100 and ImageNet-R into 10 tasks to create Split CIFAR-100 (i.e., 10 classes per task) and Split ImageNet-R (i.e., 20 classes per task), respectively [l2p, dualp, codap]. For PlantDisease, we drop 3 plant disease classes with very few images and organize the remaining 35 classes into 7 tasks to create Split PlantDisease (i.e., 5 classes per task). We also present additional results with varying task lengths (e.g., 5 tasks and 20 tasks), and an additional Split CUB-200 dataset [CUB200] in Appendix˜F and Appendix˜G, respectively.
Methods.  We employ L2P [l2p], DualPrompt [dualp], CODA-Prompt [codap], HiDe-Prompt [hidep], ConvPrompt [convp], LAE [lae], and OVOR [ovor] as representative prompting methods. These methods capitalize on ImageNet pre-trained models as backbones: ViT-L (307M), ViT-B (86M), and ViT-Ti (5.8M)—ViT-Ti is one of the smallest vision transformer models to our knowledge. Detailed settings for all methods are provided in Appendix˜B.
Hardware and metrics.  We use an NVIDIA RTX 3090 to cover a wide range of memory capacities while maintaining consistent computational power. Specifically, we limit the GPU memory to emulate resource-constrained devices, ranging from 1GB (compatible with Raspberry Pi) to 16GB (compatible with NVIDIA Jetson AGX Xavier). For computational efficiency, we measure iteration time, which is linearly correlated with GPU energy usage [miro]. We adopt the final average accuracy (higher is better) and forgetting (lower is better) from various previous works [dualp, l2p, codap, hidep]. In Appendix˜M, we further evaluate REP directly on NVIDIA Jetson TX2 with 8GB of memory, although it is applicable beyond this specific device.
@@ 4.2 Ablation studies
Table 3: Effect of using (1) conventional acceleration methods, (2) intermediate layer-based prompt selection method, and (3) random drop with diverse drop ratios, compared to our proposed methods.
@@
Algorithm validation.  Conventional token merging (ToMe) [token_merging] and layer dropping (PLD) [progressive_drop] are specifically designed to accelerate the training of transformer-based models across various domains. To validate the importance of incorporating our adaptive techniques instead in CL, we first evaluate how REP performs in case AToM and ALD are replaced with ToMe and PLD, respectively. The results are presented in (1) of Table˜3. Although applying ToMe or PLD improves resource efficiency over the native L2P baseline in Table˜1, it results in an accuracy drop of 5.1% or 2.0%, respectively, compared to using our techniques.
@@
Comparison with FIM-based token merging.  We also investigate whether additional gradient information, e.g. through the Fisher Information Matrix (FIM) [ewc], could serve as an alternative to AToM. Specifically, we compare AToM with a FIM-based, gradient-informed merging strategy using L2P with ViT-L on Split ImageNet-R (10 tasks), as shown in (3) of Table˜3. Although this approach yields modest accuracy gains (0.1–0.2%), it incurs roughly a 10% increase in training time and memory usage. By contrast, AToM preserves prompt gradients without such overhead, offering a more balanced trade-off between efficiency and forgetting mitigation. This result does not point to a fundamental limitation of FIM-based merging, but rather underscores the difficulty of achieving practical efficiency improvements in its current form.
@@
AToM and ALD intensity.  Table˜4 shows the effects of varying intensities of AToM and ALD, focusing on the number of merged tokens ($n$) in AToM and the keep ratio ($\theta$) in ALD. AToM appears to maintain stable accuracy trends even as more tokens are merged. In contrast, in the case of ALD, a lower keep ratio appears to improve resource efficiency by reducing training time. However, it can markedly impair model accuracy if the ratio is too low. Overall, when used with carefully selected hyperparameters, AToM and ALD can effectively balance resource efficiency and accuracy. Notably, forgetting remains stable when using the default settings ($n$ = 8 and $\theta$ = 0.5). In Appendix˜E, we extend the above study to two additional prompt-based CL methods, including HiDe-Prompt [hidep] and ConvPrompt [convp].